OpenStudy (sugarmochi):

Find a,b and c if n^2-n=a(n-4)^2+b(n-4)+c

OpenStudy (mathmate):

Hints: Expand left side: n^2-n=0 Expand right side: an^2+(b-8a)n+(c-4b+16a) Compare coefficients of each power, so a=1, (b-8a)=-1, (c-4b+16a)=0 Solve for a, b and c.